# Orthogonal roots, quantum Hafnians, and generalized Rothe diagrams

### v1

@@ 5.3. The invariant cubic polynomial in type $E_{6}$
The sets $U=U(E_{8},8)$ and $U^{\prime}=U(E_{7},7)$ from Theorems 5.9 and 5.13 have representation-theoretic interpretations. We recall that the exceptional Lie algebras of types $E_{6}$, $E_{7}$, and $E_{8}$ have three minuscule representations between them: two $27$-dimensional representations in type $E_{6}$, $L(E_{6},\omega_{1})$ and $L(E_{6},\omega_{6})$, and a $56$-dimensional representation in type $E_{7}$, $L(E_{7},\omega_{7})$, where each $\omega_{i}$ indicates the fundamental weight that appears as the highest weight of the corresponding representation (see, for example, [5, Theorem 5.1.5]). If one restricts the adjoint representation of the simple Lie algebra ${\mathfrak{g}}(E_{8})$ to ${\mathfrak{g}}(E_{7})$, then the set $U$ corresponds to the weights of $L(E_{7},\omega_{7})$ [5, Proposition 10.5.5]. Similarly, restricting the adjoint representation of ${\mathfrak{g}}(E_{7})$ to ${\mathfrak{g}}(E_{6})$ identifies $U^{\prime}$ with the set of weights of $L(E_{6},\omega_{1})$.
The sets $U$ and $U^{\prime}$ also have natural connections to algebraic geometry: the roots in $U$ are in bijection with the 56 exceptional curves on a del Pezzo surface $\mathcal{S}$ of degree 2, and the roots in $U^{\prime}$ are in bijection with the 27 lines on a del Pezzo surface $\mathcal{S}^{\prime}$ of degree 3 [5, Sections 9.3 and 10.1]. Under these bijections, two roots are orthogonal if and only if the intersection number between their corresponding curves is equal to $1$, so the triples of $\Omega_{U^{\prime}}$ are in a bijective correspondence with the tritangent planes of $\mathcal{S}^{\prime}$. The quadruples of $\Omega_{U}$ corresponds bijectively to the what are sometimes called “tetrads” on $\mathcal{S}$; see, for example, [4, Proposition 5.4]. Theorems 5.9 and 5.13 imply that the tritangent planes and tetrads carry natural partial orders induced by the quasiparabolic orders of $\Omega_{U^{\prime}}$ and $\Omega_{U}$, respectively. The existence of these induced orders are new to the best of our knowledge.
In the rest of this subsection, we discuss two applications of the quasiparabolic set $\Omega_{U^{\prime}}$ arising from the QP-triple $T(E_{7},7)$. First, we give explicit descriptions of the generalized Rothe diagrams associated to the elements $\Omega_{U^{\prime}}$ in Theorem 5.17. These descriptions are also new to our knowledge, and we expect they may have applications to Schubert varieties or del Pezzo surfaces. Secondly, we show in Theorem 5.19 that the Pfaffian of $\Omega_{U^{\prime}}$ agrees with the well-known invariant cubic polynomial associated to the minuscule representation $L(E_{6},\omega_{1})$. To state the theorems, we will use the following labelling conventions for the roots of $U$ adapted from the context of del Pezzo surfaces. We note that even though Theorems 5.17 and 5.19 both focus on properties of $\Omega_{U^{\prime}}$, it it more convenient to explain these conventions via the set $U$.
@@
In this realization, the $126$ roots of type $E_{7}$ consist of the $56$ coordinate permutations of $4(\varepsilon_{1}-\varepsilon_{2})$ and the $70$ coordinate permutations of $\alpha^{\prime}_{7}$; a root is positive if and only if its rightmost nonzero coefficient is negative. More details about this realization can be found in [5, Example 8.1.7].
@@
where $\tau(i)$ refers to the same vertex labeled by $i$ in the Dynkin diagram in the convention of Figure 2 (d). It is known that the map $f$ is a bijection between $U$ and the set of vectors $\{\pm v_{i,j}:1\leq i<j\leq 8\}$, where $v_{i,j}=4(\varepsilon_{i}+\varepsilon_{j})-\sum_{k=1}^{8}\varepsilon_{k}$ for all $i,j$ ([5, Example 8.1.7]). We will write $X_{ij}$ for $v_{i,j}$ and $Y_{ij}$ for $v_{i,j}$. We will also identify each vector $X_{ij}$ or $Y_{ij}$ with the corresponding root in $U$ from now on.
@@
    $\text{\rm Res}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ki}:j<k\leq 6\text{\ and\ }k\neq i\}.$
@@
    If $M:=\{\{i_{1},j_{1}\},\{i_{2},j_{2}\},\{i_{3},j_{3}\}\}$ is a perfect matching of $[6]$, then the triple $c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}$ has level $(\ell(\tau)+9)/2$, where $\tau$ is the fixed point free involution associated to $M$ and $\ell(\tau)$ is the length of $\tau$. We have
    $\text{\rm Res}_{U^{\prime}}(c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}})=\{a_{i}:1\leq i\leq 6\}\ \cup\ \{c_{pq}:\tau(p)<q\text{\ and\ }\tau(q)<p\}.$
@@
$\text{\rm Res}_{U^{\prime}}(T)=\{\beta\in U^{\prime}:s_{\gamma}(\beta)>0\text{ for all }\gamma\in T\}=\bigcap_{\gamma\in T}(U^{\prime}\setminus\{\beta\in U^{\prime}:s_{\gamma}(\beta)<0\}),$
where each set appearing in the intersection can be calculated using Lemma 5.16. The descriptions of the residues follow from the lemma by straightforward computations. Taking cardinalities of the generalized Rothe diagrams yields the levels of the triples in (i) and (ii), and the claim about the level in (iii) follows from Theorem 3.10 (v). ∎
@@
Figure 6 depicts the generalized Rothe diagrams of the elements $T_{1}=a_{4}c_{24}b_{2}$ and $T_{2}=c_{13}c_{25}c_{46}$, in the following way. First, we identify the 27 roots of $U^{\prime}$ with the 27 square cells in the array shown in Figure 6 (a) as follows: identify the cells in the top row with $a_{1},a_{2},\ldots,$ and $a_{6}$ from left to right, the cells in the rightmost column with $b_{1},b_{2},\ldots,$ and $b_{6}$ from top to bottom, and the unique cell in the $(a_{j},b_{i})$-position (i.e., the cell in the same row as $b_{i}$ and same column as $a_{j}$) with $c_{ij}=c_{ji}$. If we represent each $T_{i}$ by marking the elements of $T$ with solid black dots, then the residues of $T_{i}$ are the cells marked with a star. For an element of the form $T=a_{i}c_{ij}b_{j}$ where $j<i$, such as $T_{1}$, Theorem 5.17 (ii) implies that $\text{\rm Res}_{U^{\prime}}(T)$ consists of the cells to the left of $a_{i}$, or below $c_{ji}$, or to the left of $b_{i}$ (which is below $b_{j}$ and not in $T$). For an element of the form $T=c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}\in\Omega_{U^{\prime}}$, such as $T_{2}$, one can show using Theorem 5.17 (iii) that $\text{\rm Res}_{U^{\prime}}(T)$ consists of all the six roots in the top row and the roots $c_{ij}$ in the 27-cell grid that are neither attacked from the right nor from below by any of the roots in $T$ or their mirror images across the diagonal consisting of the $(a_{k},b_{k})$-positions for $k\in\{1,2,\ldots,6\}$. We have shaded the cell $b_{i}=b_{4}$ in Figure 6 (a) and the aforementioned mirror images in Figure 6 (b).
@@
Theorem 5.17 allows us to compute the quantum Hafnian $\text{\rm QHf}(U^{\prime})$ of $\Omega_{U^{\prime}}$ explicitly. We have
@@
We now show that the Pfaffian $\text{\rm Pf}({U^{\prime}})=\text{\rm QHf}(U^{\prime},-1)$ associated to the QP-triple $T(E_{7},7)$ agrees with the invariant cubic polynomial associated with the Lie algebra (or Chevalley group) of type $E_{6}$ as constructed by Vavilov–Luzgarev–Pevzner [17]. This provides a nonrecursive combinatorial interpretation of the coefficients of the polynomial in terms of the parity of the cardinalities of the associated generalized Rothe diagrams.
@@
The Pfaffian $\text{\rm Pf}(U^{\prime})$ associated to the QP-triple $T(E_{7},7)$ is the invariant cubic polynomial $F$ constructed in Vavilov–Luzgarev–Pevzner [17, §12].
@@
In [17, §12], Vavilov–Luzgarev–Pevzner define the minimal $W(E_{6})$-element $\{b_{6},c_{16},a_{1}\}\in\Omega_{U^{\prime}}$ from Example 5.15 to be the “distinguished triple”, which they denote as $x_{0}=(\lambda_{0},\mu_{0},\nu_{0})$. They define the invariant cubic polynomial $F$ to be the linear combination of the $45$ elements in the $W(E_{6})$-orbit of $x_{0}$ in which the coefficient of any monomial $x_{T}$ is $(-1)^{\ell(w)}$, where $w$ is an element of minimum length such that $w(x_{0})=x_{T}$. As recalled in Section 3, the length of such an element is exactly the level of $x_{T}$ by general properties of quasiparabolic sets, so we have $F=\text{\rm Pf}(\Omega_{U^{\prime}})$. ∎
@@
    Vavilov–Luzgarev–Pevzner note in [17, §12] that if we apply a simple reflection $s_{i}$ on one of the monomials $x_{T}$ in the invariant cubic polynomial, then either $s_{i}$ fixes $x_{T}$, or the coefficients of $x_{T}$ and $s_{i}(x_{T})$ are the negatives of each other. This property also characterizes $\text{\rm Pf}(U^{\prime})$ and provides an alternative proof of Theorem 5.19.
@@
    The cubic polynomial $F$ is written down explicitly in [17, Table 11]. This makes it possible, after translating notation, to verify Theorem 5.19 by hand.
@@ 6. Concluding remarks
The original Rothe diagrams and the Rothe diagrams of fixed point free involutions are useful in the study of Schubert varieties. It would be interesting to know if this also applies to the generalized Rothe diagrams of Theorem 5.17, or to the generalized Rothe diagrams associated to the sets of all positive roots of types $E_{7}$ and $E_{8}$.
@@
In the setting of Section 5.3, the 56 vectors $X_{ij}$ and $Y_{ij}$ arising from the set $U$ can be grouped into 28 pairs $i,j]:=\{X_{ij},Y_{ij}\}$. These pairs correspond naturally to the 28 bitangents to a plane quartic curve, and the 630 quadruples of $\Omega_{U}$ correspond to the 315 “tetrads of bitangents” of the quartic curve [4, Remark 5.5]. It is possible to modify the Pfaffian of $\Omega_{U}$ in a precise way to produce an invariant quartic polynomial known as Dickson’s form (see [4, Definition 5.3] and [2]), and this fact may be viewed as an analogue of Theorem 5.19 for $\Omega_{U}$. We have chosen not to describe the modification here because it requires elaborate rescalings of coefficients for which we have no conceptual explanation. However, we note that the study of geometric objects such as the 630 tetrads corresponding to $\Omega_{U}$ and the 315 tetrads of bitangents has a long history dating back to the works of Cayley and Salmon (see, for example, [4, Remark 5.5] and [15, Section 262]). It would be interesting to know if the set $\Omega_{U}$ has further implications for these classical objects.
@@
For any Coxeter system $(W,S)$ and any subset $I\subseteq S$, the left cosets of the parabolic subgroup $W_{I}\leq W$ form a quasiparabolic $W$-set whose level function $\ell_{I}$ measures the lengths of shortest coset representatives, and using $\ell_{I}$ one can deform the natural action of $W$ on the set $W/W_{I}$ to construct a module for the Iwahori–Hecke algebra $\mathcal{H}$ of $W$. The defining properties of a quasiparabolic set $X$ generalize the properties of $W/W_{I}$ and allows a similar deformation of the $W$-action on $X$ [14, Theorem 7.1]. It would be interesting to use $k$-roots to study the $\mathcal{H}$-modules obtained via the deformation of the quasiparabolic sets $X=\Omega_{U}$ constructed in this paper.
Remark 6.4.
The examples of quasiparabolic sets given in this paper all share some additional structural features that are not required by the definitions or needed for the proofs, which suggests the possible existence of an undiscovered underlying structure. For example, we have no general conceptual explanation for why the Poincaré polynomial $\sum_{R\in\Omega_{U}}q^{\rho_{U}(R)}$ always factorizes as the product of quantum integers. Furthermore, the largest such integer always turns out to be equal to $|U|/\kappa$, where $\kappa$ is the size of a maximal orthogonal subset of $U$. It is often (but not always) the case that $U$ has a partition whose blocks are elements of $\Omega_{U}$, which gives a combinatorial proof that $\kappa$ divides $|U|$. In particular, such a partition always exists for the quasiparabolic sets of the form $T(E_{n},p)$, except in the cases $T(E_{6},2)$, $T(E_{7},2)$, and $T(E_{8},5)$.
Remark 6.5.
One can create a directed acyclic graph (DAG) from a QP-triple $(W,I,U)$ by taking $U$ to be the set of vertices, and installing an edge from $u_{1}$ to $u_{2}$ if $u_{2}-u_{1}$ is a positive root of $W_{I}$. If $T\in\Omega_{U}$ satisfies $\rho(T)=0$, then $T$ is a a kernel of the DAG, meaning that (a) no two elements of $T$ are joined by an edge, and (b) for all $v\in U\backslash T$, there is an edge from $v$ to some element $u\in U$.
We can use the DAG to create a two-player game, as in Example 4.9 (ii), so that the kernel is the kernel in the sense of game theory [18, §51]. The vertices of the graph are the game positions, the directed edges are the legal moves, and playing a move that leads into the kernel constitutes a winning strategy.

### v2

@@ 5.3. The invariant cubic polynomial in type $E_{6}$
The sets $U=U(E_{8},8)$ and $U^{\prime}=U(E_{7},7)$ from Theorems 5.9 and 5.13 have representation-theoretic interpretations. We recall that the exceptional Lie algebras of types $E_{6}$, $E_{7}$, and $E_{8}$ have three minuscule representations between them: two $27$-dimensional representations in type $E_{6}$, $L(E_{6},\omega_{1})$ and $L(E_{6},\omega_{6})$, and a $56$-dimensional representation in type $E_{7}$, $L(E_{7},\omega_{7})$, where each $\omega_{i}$ indicates the fundamental weight that appears as the highest weight of the corresponding representation (see, for example, [9, Theorem 5.1.5]). If one restricts the adjoint representation of the simple Lie algebra ${\mathfrak{g}}(E_{8})$ to ${\mathfrak{g}}(E_{7})$, then the set $U$ corresponds to the weights of $L(E_{7},\omega_{7})$ [9, Proposition 10.5.5]. Similarly, restricting the adjoint representation of ${\mathfrak{g}}(E_{7})$ to ${\mathfrak{g}}(E_{6})$ identifies $U^{\prime}$ with the set of weights of $L(E_{6},\omega_{1})$.
The sets $U$ and $U^{\prime}$ also have natural connections to algebraic geometry: the roots in $U$ are in bijection with the 56 exceptional curves on a del Pezzo surface $\mathcal{S}$ of degree 2, and the roots in $U^{\prime}$ are in bijection with the 27 lines on a del Pezzo surface $\mathcal{S}^{\prime}$ of degree 3 [9, Sections 9.3 and 10.1]. Under these bijections, two roots are orthogonal if and only if the intersection number between their corresponding curves is equal to $1$, so the triples of $\Omega_{U^{\prime}}$ are in a bijective correspondence with the tritangent planes of $\mathcal{S}^{\prime}$. The quadruples of $\Omega_{U}$ correspond bijectively to the what are sometimes called “tetrads” on $\mathcal{S}$; see, for example, [6, Proposition 5.4]. Theorems 5.9 and 5.13 imply that the tritangent planes and tetrads carry natural partial orders induced by the quasiparabolic orders of $\Omega_{U^{\prime}}$ and $\Omega_{U}$, respectively. The existence of these induced orders is new to the best of our knowledge.
In the rest of this subsection, we discuss two applications of the quasiparabolic set $\Omega_{U^{\prime}}$ arising from the QP-triple $T(E_{7},7)$. First, we give explicit descriptions of the generalized Rothe diagrams associated to the elements $\Omega_{U^{\prime}}$ in Theorem 5.17. We expect that these diagrams, which are also new to our knowledge, may have applications to Schubert varieties or del Pezzo surfaces. Secondly, we show in Theorem 5.19 that the generalized Pfaffian of $\Omega_{U^{\prime}}$ agrees with the well-known invariant cubic polynomial associated to the minuscule representation $L(E_{6},\omega_{1})$. To state the theorems, we will use the following labelling conventions for the roots of $U$ adapted from the context of del Pezzo surfaces. We note that even though Theorems 5.17 and 5.19 both focus on properties of $\Omega_{U^{\prime}}$, it it more convenient to explain these conventions via the set $U$.
@@
In this realization, the $126$ roots of type $E_{7}$ consist of the $56$ coordinate permutations of $4(\varepsilon_{1}-\varepsilon_{2})$ and the $70$ coordinate permutations of $\alpha^{\prime}_{7}$; a root is positive if and only if its rightmost nonzero coefficient is negative. More details about this realization can be found in [9, Example 8.1.7].
@@
where $\tau(i)$ refers to the same vertex labeled by $i$ in the Dynkin diagram in the convention of Figure 2 (d). It is known that the map $f$ is a bijection between $U$ and the set of vectors $\{\pm v_{i,j}:1\leq i<j\leq 8\}$, where $v_{i,j}=4(\varepsilon_{i}+\varepsilon_{j})-\sum_{k=1}^{8}\varepsilon_{k}$ for all $i,j$ ([9, Example 8.1.7]). We will write $X_{ij}$ for $v_{i,j}$ and $Y_{ij}$ for $v_{i,j}$. We will also identify each vector $X_{ij}$ or $Y_{ij}$ with the corresponding root in $U$ from now on.
@@
    $\mathscr{D}_{U^{\prime}}(a_{i}c_{ij}b_{j})=\{a_{h}:1\leq h<i\}\ \cup\ \{c_{ki}:j<k\leq 6\text{\ and\ }k\neq i\}.$
@@
    If $M:=\{\{i_{1},j_{1}\},\{i_{2},j_{2}\},\{i_{3},j_{3}\}\}$ is a perfect matching of $[6]$, then the triple $c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}$ has level $(\ell(\tau)+9)/2$, where $\tau$ is the fixed-point-free involution associated to $M$ and $\ell(\tau)$ is the length of $\tau$. We have
    $\mathscr{D}_{U^{\prime}}(c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}})=\{a_{i}:1\leq i\leq 6\}\ \cup\ \{c_{pq}:\tau(p)<q\text{\ and\ }\tau(q)<p\}.$
@@
$\mathscr{D}_{U^{\prime}}(T)=\{\beta\in U^{\prime}:s_{\gamma}(\beta)>0\text{ for all }\gamma\in T\}=\bigcap_{\gamma\in T}(U^{\prime}\setminus\{\beta\in U^{\prime}:s_{\gamma}(\beta)<0\}),$
where each set appearing in the intersection can be calculated using Lemma 5.16. The descriptions of the generalized Rothe diagrams follow from the lemma by straightforward computations. Taking cardinalities of the generalized Rothe diagrams yields the levels of the triples in (i) and (ii), and the claim about the level in (iii) follows from Theorem 3.11 (v). ∎
@@
Figure 6 depicts the generalized Rothe diagrams of the elements $T_{1}=a_{4}c_{24}b_{2}$ and $T_{2}=c_{13}c_{25}c_{46}$, in the following way. First, we identify the 27 roots of $U^{\prime}$ with the 27 square cells in the array shown in Figure 6 (a) as follows: identify the cells in the top row with $a_{1},a_{2},\ldots,$ and $a_{6}$ from left to right, the cells in the rightmost column with $b_{1},b_{2},\ldots,$ and $b_{6}$ from top to bottom, and the unique cell in the $(a_{j},b_{i})$-position (i.e., the cell in the same row as $b_{i}$ and same column as $a_{j}$) with $c_{ij}=c_{ji}$. If we represent each $T_{i}$ by marking the elements of $T$ with solid black dots, then the roots dominating $T_{i}$ are the cells marked with a star. For an element of the form $T=a_{i}c_{ij}b_{j}$ where $j<i$, such as $T_{1}$, Theorem 5.17 (ii) implies that $\mathscr{D}_{U^{\prime}}(T)$ consists of the cells to the left of $a_{i}$, or below $c_{ji}$, or to the left of $b_{i}$ (which is below $b_{j}$ and not in $T$). For an element of the form $T=c_{i_{1}j_{1}}c_{i_{2}j_{2}}c_{i_{3}j_{3}}\in\Omega_{U^{\prime}}$, such as $T_{2}$, one can show using Theorem 5.17 (iii) that $\mathscr{D}_{U^{\prime}}(T)$ consists of all the six roots in the top row and the roots $c_{ij}$ in the 27-cell grid that are neither attacked from the right nor from below by any of the roots in $T$ or their mirror images across the diagonal consisting of the $(a_{k},b_{k})$-positions for $k\in\{1,2,\ldots,6\}$. We have shaded the cell $b_{i}=b_{4}$ in Figure 6 (a) and the aforementioned mirror images in Figure 6 (b).
@@
Theorem 5.17 allows us to compute $\text{\rm QHf}(U^{\prime})$ explicitly. We have
@@
We now show that the generalized Pfaffian $\text{\rm Pf}({U^{\prime}})=\text{\rm QHf}(U^{\prime},-1)$ associated to the QP-triple $T(E_{7},7)$ agrees with the invariant cubic polynomial associated with the Lie algebra (or Chevalley group) of type $E_{6}$ as constructed by Vavilov–Luzgarev–Pevzner [22]. This provides a nonrecursive combinatorial interpretation of the coefficients of the polynomial in terms of the parity of the cardinalities of the associated generalized Rothe diagrams.
@@
The generalized Pfaffian $\text{\rm Pf}(U^{\prime})$ associated to the QP-triple $T(E_{7},7)$ is the invariant cubic polynomial $F$ constructed in Vavilov–Luzgarev–Pevzner [22, §12].
@@
In [22, §12], Vavilov–Luzgarev–Pevzner define the minimal $W(E_{6})$-element $\{b_{6},c_{16},a_{1}\}\in\Omega_{U^{\prime}}$ from Example 5.15 to be the “distinguished triple”, which they denote as $x_{0}=(\lambda_{0},\mu_{0},\nu_{0})$. They define the invariant cubic polynomial $F$ to be the linear combination of the $45$ elements in the $W(E_{6})$-orbit of $x_{0}$ in which the coefficient of any monomial $x_{T}$ is $(-1)^{\ell(w)}$, where $w$ is an element of minimum length such that $w(x_{0})=x_{T}$. As recalled in Section 3, the length of such an element is exactly the level of $x_{T}$ by general properties of quasiparabolic sets, so we have $F=\text{\rm Pf}(\Omega_{U^{\prime}})$. ∎
@@
    Vavilov–Luzgarev–Pevzner note in [22, §12] that if we apply a simple reflection $s_{i}$ to one of the monomials $x_{T}$ in the invariant cubic polynomial, then either $s_{i}$ fixes $x_{T}$, or the coefficients of $x_{T}$ and $s_{i}(x_{T})$ are the negatives of each other. This property also characterizes $\text{\rm Pf}(U^{\prime})$ and provides an alternative proof of Theorem 5.19.
@@
    The cubic polynomial $F$ is written down explicitly in [22, Table 11]. This makes it possible, after translating notation, to verify Theorem 5.19 by hand.
@@
Remark 5.21.
It is possible to establish an analogue of Theorem 5.19 for $\Omega_{U}$ as follows. The 56 vectors $X_{ij}$ and $Y_{ij}$ arising from the set $U$ can be grouped into 28 pairs $i,j]:=\{X_{ij},Y_{ij}\}$. These pairs correspond naturally to the 28 bitangents to a plane quartic curve, and the 630 quadruples of $\Omega_{U}$ correspond to the 315 “tetrads of bitangents” of the quartic curve [6, Remark 5.5]. One can modify the generalized Pfaffian of $\Omega_{U}$ in a precise way to produce an invariant quartic polynomial known as Dickson’s form (see [6, Definition 5.3] and [4]), thus providing the promised analogue of Theorem 5.19 for $\Omega_{U}$. We have chosen not to describe the modification here because it requires elaborate rescalings of coefficients for which we have no conceptual explanation. However, we note that the study of geometric objects such as the 630 tetrads corresponding to $\Omega_{U}$ and the 315 tetrads of bitangents has a long history dating back to the works of Cayley and Salmon (see, for example, [6, Remark 5.5] and [20, Section 262]). It would be interesting to know if the set $\Omega_{U}$ has further implications for these classical objects.
@@ 6. Concluding remarks
The original Rothe diagrams and the Rothe diagrams of fixed-point-free involutions are useful in the study of Schubert varieties. It would be interesting to know if this also applies to the generalized Rothe diagrams of Theorem 5.17, or to the generalized Rothe diagrams associated to the sets of all positive roots of types $E_{7}$ and $E_{8}$.
@@
For any Coxeter system $(W,S)$ and any subset $I\subseteq S$, the left cosets of the parabolic subgroup $W_{I}\leq W$ form a quasiparabolic $W$-set whose level function $\ell_{I}$ measures the lengths of shortest coset representatives, and using $\ell_{I}$ one can deform the natural action of $W$ on the set $W/W_{I}$ to construct a module for the Iwahori–Hecke algebra $\mathcal{H}$ of $W$. The defining properties of a quasiparabolic set $X$ generalize the properties of $W/W_{I}$ and allows a similar deformation of the $W$-action on $X$ [19, Theorem 7.1]. It would be interesting to use $k$-roots to study the $\mathcal{H}$-modules obtained via the deformation of the quasiparabolic sets $X=\Omega_{U}$ constructed in this paper.
@@
For all the quasiparabolic sets $\Omega_{U}$ constructed in this paper, the Poincaré polynomial factors as a multiplicity-free product of quantum integers. The significance of the corresponding integers, i.e., the degrees of $U$, may deserve further study. In particular, it would be interesting to know why the degrees can be computed via the matroid associated to $\Omega_{U}$ as explained in Section 4.3. For the QP-triples $\Omega_{U}$ from Section 4.2, it is also natural to wonder whether the degrees of $U$ have interpretations in terms of the corresponding symmetric spaces and symmetric pairs.